Question 20 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes. With specialization, Spain produces only grapes-75 tons-and France produces 75 ships. The terms of trade are 1 ship for 1 ton of grapes, and 30 ships are traded for 30 tons of grapes. Are the countries better off or worse off with trade? With specialization and trade, how many additional tons nsume? Provide your answer as a number rounded to two decimal places Do not include any symbols, such as s, r, %, or , in your Use - to indicate negative amounts. France Spain Grapes (in tons) 50 40 30 Ships Grapes (in tons) 75 Ships 10 20 30 40 50 30 45 45 30 15 75 Your Answer:
Question 21 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Suppose that Spaín is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes With specialization, Spain produces only grapes-75 tons-and France produces 75 ships. The terms of trade are 1 ship for 1 ton of grapes, and 30 ships are traded for 30 tons of grapes. Are the countries better off or worse off with trade? With specialization and trade, how many additional ships (if any) can Spain consume? Provide your answer as a number rounded to two decimal places. Use - to indicate negative amounts. Do not include any symbols, such as S, , %, or , in your answer. Spain France Grapes (in tons) 50 40 30 20 10 Ships Grapes (in tons) 75 60 45 30 15 Shi 0 10 20 30 40 50 60 75 Your Answer Answer
Question 22 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes. With specialization, Spain produces only grapes-75 tons-and France produces 75 ships. The terms of trade are 1 ship for 1 ton of grapes, and 30 ships are traded for 30 tons of grapes. Are the countries better off or worse off with trade? With specialization and trade, how many additional ships (if any) can France consume? Provide your answer as a number rounded to two decimal places. Use - to indicate negative amounts. Do not include any symbols, such as s, , %, or , in your answer. Spain France Ships Grapes (in tons) 75 60 45 30 15 Grapes (in tons) 50 40 30 20 10 Ships 10 20 30 40 50 15 30 45 75 Your Answer
Question 23 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price i both countries. The table below shows the combinations of the two goods that each cournt produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes With specialization, Spain produces only grapes-75 tons-and France produces 75 ships. The terms of trade are 1 ship for 1 ton of grapes, and 30 ships are traded for 30 tons of grapes. Are the countries better off or worse off with trade? With specialization and trade, how many additional tons of grapes (if any) can France consume? Provide your answer as a number rounded to two decimal places. Use - to indicate negative amounts. Do not include any symbols, such as s, , %, or , in your answer. France Spain Grapes (in tons) 50 40 30 20 10 Grapes (in tons) Ships Shi 15 30 10 20 30 40 50 60 45 30 15 75 Your Answer: Answer
Answer #1

20) Without trade, Spain consumers 30 tons of grapes and after trade Spain consumes 75-30= 45 tons of grapes.

Additional tons of grapes consumed by Spain= 45-30= 15

Answer. 15

21) Spain consumes the same amount of ships before and after trade.

Additional number of ships= 0

Answer. 0

22) additional number of ships France can consume= 75-30= 45

Answer. 45

23) additional tons of grapes France can consume= 0

Answer. 0